OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? x − y = 8 −x − y = −7

OpenStudy (anonymous):

A.8 − y B. −8 − y C. 8 + y D. −8 + y

OpenStudy (nurali):

x − y = 8 x=8+y

OpenStudy (anonymous):

x-y=8 x=8+y So C

OpenStudy (anonymous):

Thank you so much!

OpenStudy (nurali):

Anytime.